# SVD-LLM: Truncation-aware Singular Value Decomposition for Large Language Model Compression

### v2

@@ 3.1 Truncation-Aware Data Whitening
Theorem 3.2.
If the whitening matrix $S$ satisfies $SS^{T}=XX^{T}$, the compression loss $L_{i}$ equals to $\tilde{\sigma_{i}}$.
Proof.
Since $\tilde{X}=S^{-1}X$ and $SS^{T}=XX^{T}$, we can further infer Section 3.1 to get:
$L_{i}$ $=||\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}||_{F}=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\tilde{X}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}S^{-1}XX^{T}(S^{-1})^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}S^{-1}SS^{T}(S^{T})^{-1}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}$
$=\tilde{\sigma}_{i}~{}trace(\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{v}_{i}\tilde{u}_{i}^{T})^{\frac{1}{2}}=\tilde{\sigma}_{i}$
Since $S$ is the Cholesky decomposition of $XX^{T}$ in our algorithm, which satisfies $SS^{T}=XX^{T}$, the loss $L_{i}$ of truncating $\tilde{\sigma_{i}}$ equals to the singular value $\tilde{\sigma_{i}}$ itself. ∎
Corollary 3.3.
If the whitening matrix $S$ satisfies $SS^{T}=XX^{T}$, truncating the smallest singular values leads to the lowest loss $L$ compared to truncating others.
Proof.
Suppose we truncate $\tilde{\sigma}_{m+1},\tilde{\sigma}_{m+2},\tilde{\sigma}_{m+3},...,\tilde{\sigma}_{r}$ in $\tilde{\Sigma}$ for compression, we calculate the square of the loss $L$:
$L^{2}=$ $\left|\left|\sum_{i=m+1}^{r}\tilde{\sigma}_{i}\tilde{u}_{i}\tilde{v}_{i}^{T}\tilde{X}\right|\right|^{2}_{F}$
$=$ $\sum_{j=m+1}^{r}\sum_{i=m+1}^{r}\tilde{\sigma}_{i}\tilde{\sigma}_{j}~{}trace(\tilde{u_{i}}\tilde{v_{i}}^{T}\tilde{X}\tilde{X}^{T}\tilde{v_{j}}\tilde{u_{j}}^{T})$
$=$ $\sum_{i=m+1}^{r}\tilde{\sigma}_{i}^{2}~{}trace(\tilde{u_{i}}\tilde{v_{i}}^{T}\tilde{X}\tilde{X}^{T}\tilde{v_{i}}\tilde{u_{i}}^{T})$
$=$ $\sum_{i=m+1}^{r}(L_{i})^{2}=\sum_{i=1}^{k}(\tilde{\sigma}_{i})^{2}$
The squared loss $L^{2}$ is equal to the sum of the squared singular values. Therefore, truncating the smallest singular values $\tilde{\sigma}_{i}$ achieves the lowest compression loss. ∎
3.2 Layer-Wise Closed-Form Update
Motivation: With the increase of model compression ratio, the weight matrices in the compressed LLM generally lose their ability to adapt to the new activation $X^{\prime}$ derived from the compressed LLM and thus requires the external update for the adaptation during the model compression process. However, existing SVD-based LLM compression techniques do not include this update process, leading to less competitive performance at high compression ratios.
Table 1: Perplexity and Average Adaptation Loss of LLaMA-7B compressed by ASVD on WikiText-2. The ratio of increase in both perplexity and average adaptation loss from the previous compression ratio to the current one is marked in red.
Metric Compression Ratio
20% 30% 40% 50%
Perplexity 11.14 51 (x4.6) 1407 (x28) 15358 (x11)
Average Adaptation Loss 0.83 2.98 (x3.6) 9.63 (x3.2) 44.42 (x4.6)
As listed in Table 1, the perplexity of LLaMA-7B compressed by ASVD increases about 28 times when the compression ratio increases from 30% to 40%. Meanwhile, there is a significant increase in the average of the adaptation loss computed by $L_{a}=||WX^{\prime}-W^{\prime}X^{\prime}||_{F}$, which demonstrates the significant role of minimizing $L_{a}$, especially under high compression ratios.
Key Design: The key idea of SVD-LLM is to incorporate a layer-wise closed-form model parameter update strategy to update the remaining weights after SVD truncation that minimizes the adaptation loss $L_{a}$:
$O_{2}$ $=\min(||WX^{\prime}-W^{\prime}X^{\prime}||_{F})$ (4)
SVD-LLM compresses and updates all the local weights in one decoder layer at a time. In order to perform the update without destroying the low-rank structure of the compressed weights, SVD-LLM only updates the matrix $\tilde{U}$ while keeping $\tilde{\Sigma}_{1}$ (the singular value matrix after truncation) and $\tilde{V}$ fixed:
$\tilde{U^{\prime}}=\arg\min_{\tilde{U^{\prime}}}||WX^{\prime}-\tilde{U^{\prime}}\tilde{\Sigma}_{1}\tilde{V}^{T}X^{\prime}||_{F}$ (5)
To compensate for the error introduced by truncation, we adapt $\tilde{U}$ to its closed-form solution $\tilde{U^{\prime}}=WX^{\prime}X^{\prime T}\tilde{V}\tilde{\Sigma}_{1}(\tilde{\Sigma}_{1}\tilde{V}^{T}X^{\prime}X^{\prime T}\tilde{V}\tilde{\Sigma}_{1})^{-1}$ with the same calibration data used for whitening.
Figure 3: Layer-wise closed-form update in SVD-LLM.
Figure 3 illustrates the overall process. To perform the update in layer $i$, we extract the output $x^{\prime}_{i-1}$ from the previous updated layer $i-1$ as the new input activation, then apply it to every original weight matrix $W$ and its corresponding decomposed matrices $\tilde{\Sigma}_{1}$ and $\tilde{V}$ in layer $i$ to update $\tilde{U}$ based on Equation 5.
After updating $\tilde{U}$ with $\tilde{U^{\prime}}$, we combine and save the truncated matrices $\tilde{U^{\prime}},\tilde{\Sigma}_{1},\tilde{V}$ into two low-rank weight matrices $W_{u}=\tilde{U^{\prime}}(\tilde{\Sigma}_{1})^{1/2}$ and $W_{v}=\tilde{V}(\tilde{\Sigma}_{1})^{1/2}$. Therefore, for the new input activation $X^{\prime}$, its forward pass takes the form $Y=W_{u}W_{v}^{T}X^{\prime}$ in compressed LLM.
@@ 4 Experiments
4.1 Experimental Setups
Baselines. We compare SVD-LLM against three baselines including vanilla SVD as well as state-of-the-art SVD-based LLM compression methods FWSVD (Hsu et al., 2022) and ASVD (Yuan et al., 2023).
Models and Datasets. To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on seven models from three different LLM families (LLaMA-7B, 13B, 30B, 65B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), Vicuna-7B (Chiang et al., 2023) and Mistral-7B (Jiang et al., 2023) and $11$ datasets including three language modeling datasets (WikiTex-2 (Merity et al., 2017), PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020)) and eight classification datasets (SST-5 (Socher et al., 2013), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), COLA (Warstadt et al., 2019), SNLI (Bowman et al., 2015), ARC-e, ARC-c, and ARC-b (Clark et al., 2018)) in zero-shot setting following previous experimental design (Ma et al., 2023).
Implementation Details. To ensure a fair comparison, we followed ASVD (Yuan et al., 2023) to use the calibration set of 256 samples from WikiText-2 for whitening. This calibration set is also used for layer-wise closed-form update. Since layer-wise closed-form update is intended to mitigate accuracy drop under higher compression ratios, we only apply it when compression ratios are at 40% and above. All of our experiments are conducted on Nvidia A100 GPUs.
@@ 4.2 Overall Performance
We evaluate the overall performance of SVD-LLM from five aspects: (1) performance on different datasets, (2) performance under high model compression ratios, (3) performance on different LLMs, (4) performance on LLMs with larger scales, and (5) performance with LoRA fine-tuning.
Performance on Different Datasets.
Table 2: Zero-shot performance comparison of LLaMA-7B compressed by SVD-LLM and baselines on three language modeling datasets (measured by perplexity ($\downarrow$)) and eight classification datasets (measured by both individual and average accuracy ($\uparrow$)). The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
@@
0% Original 5.68 8.35 7.34 31.54 34.43 66.49 69.13 34.26 26.63 29.49 25.75 39.72
20% SVD 20061 20306 18800 18.91 31.54 33.74 30.87 31.70 20.46 20.00 21.93 26.14
FWSVD 1727 2152 1511 26.38 32.07 42.32 49.09 33.59 23.10 26.44 24.01 32.13
ASVD 11.14 16.55 15.93 26.38 33.50 66.49 68.17 33.09 25.40 23.39 25.64 37.76
SVD-LLM 7.94 ($\downarrow$29%) 16.22 ($\downarrow$2%) 15.84 ($\downarrow$1%) 26.92 35.47 66.55 68.84 34.41 27.16 27.46 26.22 39.13 ($\uparrow$4%)
First, we evaluate the performance of LLaMA-7B compressed by SVD-LLM and the baselines under 20% compression ratio on all $11$ datasets. Table 2 summarizes the results. As shown, SVD-LLM consistently outperforms vanilla SVD, FWSVD and ASVD across all the $11$ datasets. Specifically, on the three language modeling datasets, SVD-LLM outperforms vanilla SVD and FWSVD by a large margin; in comparison with ASVD, SVD-LLM reduces the perplexity by 29%, 2%, and 1% on WikiText-2, PTB, and C4, respectively. On the eight classification datasets, SVD-LLM performs better than the best-performing baseline and achieves 4% higher accuracy on average across the eight datasets.
@@
Performance under High Model Compression Ratios. None of the existing SVD-based LLM compression methods report their performance on model compression ratios above 20%. To bridge this gap, we compare the performance of LLaMA-7B compressed by SVD-LLM and the baselines under compression ratios from 30% to 60%. Table 3 summarizes our results. As shown, SVD-LLM consistently outperforms the three baselines across all the compression ratios. More importantly, compared to low compression ratio scenario in Table 2, SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios. Specifically, under 30% compression ratio, compared to the best-performing baseline (ASVD), SVD-LLM reduces the perplexity on WikiText-2, PTB, and C4 by 81%, 62%, and 39%, respectively; under compression ratio of 40% and above, SVD-LLM reduces the perplexity by more than 96%. These results indicate that SVD-LLM is more effective in compressing LLMs for more resource-constrained devices such as smartphones and IoT devices.
Table 3: Perplexity ($\downarrow$) of LLaMA-7B compressed by SVD-LLM and baselines under high model compression ratios on WikiText-2, PTB, and C4.
Method 30 % 40% 50% 60%
WikiText-2
SVD 13103 52489 131715 105474
FWSVD 20127 18156 24391 32194
ASVD 51 1407 15358 57057
SVD-LLM 9.56 ($\downarrow$81%) 13.11 ($\downarrow$99%) 23.97 ($\downarrow$99%) 53.74 ($\downarrow$99%)
PTB
SVD 17210 59977 87227 79905
FWSVD 11058 20990 28321 43931
ASVD 70 3292 47690 45218
SVD-LLM 26.39 ($\downarrow$62%) 63.75 ($\downarrow$98%) 150.58 ($\downarrow$99%) 438.58 ($\downarrow$99%)
C4
SVD 20871 47774 79815 106976
FWSVD 7240 12847 23104 29292
ASVD 41 1109 27925 43036
SVD-LLM 25.11 ($\downarrow$39%) 49.83 ($\downarrow$96%) 118.57 ($\downarrow$99%) 345.49 ($\downarrow$99%)
@@
Performance on Different LLMs. To examine the generability of SVD-LLM across different LLMs, we compare the performance between SVD-LLM and the baselines on four models from three different LLM families – LLaMA-7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B – under 20% compression ratio on WikiText-2. As shown in Table 4, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four LLMs. In addition, SVD-LLM exhibits more stable performance on different LLM families, especially compared to vanilla SVD and FWSVD.
@@
Table 4: Perplexity ($\downarrow$) of different LLMs under 20% compression ratio on WikiText-2. The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
Method LLaMA-7B LLaMA 2-7B Mistral-7B Vicuna-7B
SVD 20061 18192 159627 18644
FWSVD 1630 2360 6357 2758
ASVD 11.14 10.10 13.72 16.23
SVD-LLM 7.94 ($\downarrow$29%) 8.50 ($\downarrow$16%) 10.21 ($\downarrow$26%) 6.78 ($\downarrow$58%)
Performance on LLMs with Larger Scales. To examine the generability of SVD-LLM on LLMs across different scales, we compare the performance between SVD-LLM and the baselines on LLaMA series at four different scales – 7B, 13B, 30B, and 65B – under 20% compression ratio on WikiText-2. As shown in Table 5, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four model sizes. Moreover, both FWSVD and ASVD demand excessive memory resources, causing out of memory (OOM) when compressing LLMs at larger scales even on an A100 GPU due to memory-intensive operations for estimating the importance of weight matrices. In contrast, SVD-LLM does not involve such estimation operations and thus avoids OOM.
@@
Performance with LoRA Fine-Tuning.
Figure 4: Perplexity of LLaMA-7B compressed by SVD-LLM and ASVD (w/ and w/o LoRA) on WikiText-2. The perplexities of ASVD at and beyond $30\%$ compression ratio are higher than $1407$ and thus are not shown in the figure.
LoRA (Hu et al., 2022) is a common fine-tuning technique for LLM. It has been applied with pruning-based LLM compression methods such as LLM-Pruner (Ma et al., 2023) to mitigate accuracy drop after pruning. LoRA can also be combined with SVD-based LLM compression methods by modifying the forward pass of a linear layer as:
$Y$ $=W_{u}^{\prime}W_{v}^{\prime}X$ (6)
To examine the performance of SVD-LLM in combination with LoRA, we follow the same configuration used in LLM-Pruner (Ma et al., 2023) to fine-tune LLaMA-7B compressed by SVD-LLM and ASVD under compression ratios from 20% to 80% with LoRA. The results are shown in Figure 4. We have three observations. (1) Comparing SVD-LLM with SVD-LLM $+$ LoRA, under the compression ratio between 20% and 50%, the accuracy enhancement brought by LoRA is limited; as the compression ratio increases, LoRA is playing a more significant role in enhancing the performance of SVD-LLM. (2) Compared with ASVD $+$ LoRA, SVD-LLM $+$ LoRA consistently achieves better accuracy across all the compression ratios. In particular, under 70% compression ratio, the perplexity of SVD-LLM $+$ LoRA is 75% lower than ASVD $+$ LoRA. (3) Lastly, even without LoRA, SVD-LLM is able to achieve perplexity comparable to ASVD $+$ LoRA, especially under the compression ratios between 20% and 60%.
@@ 4.3 Compression Speed Evaluation
Besides compression performance, we also evaluate the compression speed of SVD-LLM and the baselines. Specifically, we measured the GPU hours used for SVD-LLM and ASVD when compressing LLaMA-7B under 20% compression ratio on an A100 GPU. The results are shown in Table 6. As shown, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes, which is $32$ times faster. When breaking down the time, most of the time consumed by ASVD is dedicated to calculating the compression ratio of each weight matrix based on its estimated importance through a search process. In contrast, SVD-LLM maintains a consistent compression ratio among all the weight matrices and thus gets rid of the time-consuming search process.
Table 6: Compression speed comparison on LLaMA-7B under 20% compression ratio. The relative speedup is marked in green color inside bracket.
@@
Whitening Update Total Preprocessing Search Total
@@
4.4 Ablation Studies
We next conduct ablation studies to evaluate the separate contributions of the two key components (truncation-aware data whitening and layer-wise closed-form update) of SVD-LLM. Let SVD-LLM (W) denote the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denote the version of SVD-LLM with layer-wise closed-form update only; and SVD-LLM (W+U) denote the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update. The results are shown in Table 7. We have three observations. (1) Both SVD-LLM (W) and SVD-LLM (U) consistently outperform ASVD across all the compression ratios. Notably, when the compression ratio is at and above 40%, both variants reduce the perplexity by more than 99% compared to ASVD. (2) Under 20% and 30% compression ratios, SVD-LLM (W) achieves the lowest perplexity compared to SVD-LLM (U) and SVD-LLM (W+U). (3) Under 40%, 50% and 60% compression ratios, SVD-LLM (W+U) achieves the lowest perplexity compared to SVD-LLM (W) and SVD-LLM (U), highlighting the importance of combining both truncation-aware data whitening and layer-wise closed-form update when compression ratio goes high.
Table 7: Perplexity ($\downarrow$) of LLaMA-7B compressed under different compression ratios on WikiText-2. SVD-LLM (W) denotes the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denotes the version of SVD-LLM with layer-wise closed-form update only; SVD-LLM (W+U) denotes the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update. The relative performance gain compared to ASVD is marked in green color inside bracket.

### v3

@@ 3.2 Layer-Wise Closed-Form Update
$\begin{split}U_{i}^{\prime}&=\arg\min_{U_{i}^{\prime}}||W_{i}X^{\prime}_{i-1}-W^{\prime}_{i}X^{\prime}_{i-1}||_{F}\\ &=W_{i}X^{\prime}_{i-1}D^{T}(DD^{T})^{-1},D=\text{Trunc.}(\Sigma)_{i}V_{i}^{T}S_{i}^{-1}X^{\prime}_{i-1}\end{split}$ (7)
@@ 4 Experiments
Baselines. We compare SVD-LLM against three baselines including vanilla SVD as well as state-of-the-art SVD-based LLM compression methods FWSVD [12] and ASVD [28].
Models and Datasets. To demonstrate the generability of our method, we evaluate the performance of SVD-LLM and the baselines on eight models from three different LLM families (LLaMA-7B, 13B, 30B, 65B [23], LLaMA2-7B [24], OPT-6.7B [30], Vicuna-7B [3] and Mistral-7B [14]) and $10$ datasets including three language modeling datasets (WikiText-2 [18], PTB [17] and C4 [21]) and seven classification datasets (OpenbookQA [20], WinoGrande [22], HellaSwag [29], PIQA [2], MathQA [1], ARC-e, and ARC-c [4]) in zero-shot setting with LM-Evaluation-Harness framework [7].
Implementation Details. To ensure a fair comparison, we followed ASVD [28] to randomly select 256 samples from WikiText-2 as the calibration data. Since layer-wise closed-form update is intended to mitigate the accuracy drop under higher compression ratios, we only apply it when the compression ratios are at 40% and above. All of our experiments are conducted on Nvidia A100 GPUs.
@@ 4.1 Overall Performance
We evaluate the overall performance of SVD-LLM from four aspects: (1) performance under different compression ratios, (2) performance on different LLMs, (3) performance on LLMs with larger scales, and (4) performance with LoRA fine-tuning (See Section A.3). Some generated contents by the compressed LLM are listed in Section A.4 to provide a more straightforward comparison.
Performance under Different Compression Ratios.
Table 1: Zero-shot performance of LLaMA-7B compressed by SVD-LLM and baselines under 20% to 60% compression ratio on three language modeling datasets (measured by perplexity ($\downarrow$)) and seven common sense reasoning datasets (measured by both individual and average accuracy ($\uparrow$)). The best performance is marked in bold. The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
@@
0% Original 5.68 8.35 7.34 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.52
20% SVD 20061 20306 18800 0.14 0.27 0.51 0.26 0.21 0.53 0.21 0.31
FWSVD 1727 2152 1511 0.15 0.31 0.50 0.26 0.23 0.56 0.21 0.32
ASVD 11.14 16.55 15.93 0.25 0.53 0.64 0.41 0.27 0.68 0.24 0.43
SVD-LLM 7.94 ($\downarrow$29%) 16.22 ($\downarrow$2%) 15.84 ($\downarrow$1%) 0.22 0.58 0.63 0.43 0.29 0.69 0.24 0.44 ($\uparrow$2%)
30% SVD 13103 17210 20871 0.13 0.26 0.51 0.26 0.21 0.54 0.22 0.30
FWSVD 20127 11058 7240 0.17 0.26 0.49 0.26 0.22 0.51 0.19 0.30
ASVD 51 70 41 0.18 0.43 0.53 0.37 0.25 0.65 0.21 0.38
SVD-LLM 9.56 ($\downarrow$81%) 26.39 ($\downarrow$62%) 25.11 ($\downarrow$39%) 0.20 0.48 0.59 0.37 0.26 0.65 0.22 0.40 ($\uparrow$5%)
40% SVD 52489 59977 47774 0.15 0.26 0.52 0.26 0.22 0.53 0.20 0.30
FWSVD 18156 20990 12847 0.16 0.26 0.51 0.26 0.22 0.53 0.21 0.30
ASVD 1407 3292 1109 0.13 0.28 0.48 0.26 0.22 0.55 0.19 0.30
SVD-LLM 13.11 ($\downarrow$99%) 63.75 ($\downarrow$98%) 49.83 ($\downarrow$96%) 0.19 0.42 0.58 0.33 0.25 0.60 0.21 0.37 ($\uparrow$23%)
50% SVD 131715 87227 79815 0.16 0.26 0.50 0.26 0.23 0.52 0.19 0.30
FWSVD 24391 28321 23104 0.12 0.26 0.50 0.26 0.23 0.53 0.20 0.30
ASVD 15358 47690 27925 0.12 0.26 0.51 0.26 0.22 0.52 0.19 0.30
SVD-LLM 23.97 ($\downarrow$99%) 150.58 ($\downarrow$99%) 118.57 ($\downarrow$99%) 0.16 0.33 0.54 0.29 0.23 0.56 0.21 0.33 ($\uparrow$10%)
60% SVD 105474 79905 106976 0.16 0.26 0.50 0.26 0.22 0.52 0.21 0.30
FWSVD 32194 43931 29292 0.15 0.26 0.49 0.26 0.22 0.53 0.18 0.30
ASVD 57057 45218 43036 0.12 0.26 0.49 0.26 0.21 0.51 0.18 0.29
SVD-LLM 53.74 ($\downarrow$99%) 438.58 ($\downarrow$99%) 345.49 ($\downarrow$99%) 0.14 0.28 0.50 0.27 0.22 0.55 0.21 0.31 ($\uparrow$7%)
@@
First, we evaluate the performance of LLaMA-7B compressed by SVD-LLM and the baselines under 20% to 60% compression ratios. Section 4.1 summarizes the results on all $10$ datasets. As shown, SVD-LLM consistently outperforms vanilla SVD, FWSVD and ASVD across all of the compression ratios. More importantly, compared to the low compression ratio scenario in Section 4.1, SVD-LLM exhibits significant advantages over vanilla SVD, FWSVD, and ASVD under high compression ratios. Specifically, under 30% compression ratio, compared to the best-performing baseline (ASVD), SVD-LLM reduces the perplexity on WikiText-2, PTB, and C4 by 81%, 62%, and 39%, respectively; When the compression ratio reaches 40% and above, SVD-LLM reduces the perplexity by more than 96%. These results indicate that SVD-LLM is more effective in compressing LLMs for more resource-constrained devices such as smartphones and IoT devices. On the seven classification datasets, SVD-LLM performs better than the best-performing baseline on most of the datasets and consistently achieves at least 2% higher average accuracy across all the compression ratios.
@@
Performance on Different LLMs. To examine the generability of SVD-LLM across different LLMs, we compare the performance between SVD-LLM and the baselines on four different models, including OPT-6.7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B under 20% compression ratio on WikiText-2 and the common sense reasoning datasets (See Section A.5). The result on WikiText-2 is shown in Section 4.1, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four LLMs. In addition, SVD-LLM exhibits more stable performance on different LLM families, especially compared to vanilla SVD and FWSVD.
@@
Table 2: Perplexity ($\downarrow$) of four different LLMs including OPT-6.7B, LLaMA 2-7B, Mistral-7B, and Vicuna-7B under 20% compression ratio on WikiText-2. The relative performance gain compared to the best-performing baseline is marked in green color inside bracket.
Method OPT-6.7B LLaMA 2-7B Mistral-7B Vicuna-7B
SVD 66275 18192 159627 18644
FWSVD 14559 2360 6357 2758
ASVD 82 10.10 13.72 16.23
SVD-LLM 16.04 ($\downarrow$80%) 8.50 ($\downarrow$16%) 10.21 ($\downarrow$26%) 6.78 ($\downarrow$58%)
@@
Performance on LLMs with Larger Scales. To examine the generability of SVD-LLM on LLMs across different scales, we compare the performance between SVD-LLM and the baselines on LLaMA series at four different scales – 7B, 13B, 30B, and 65B – under 20% compression ratio on WikiText-2. As shown in Section 4.1, SVD-LLM consistently outperforms vanilla SVD, FWSVD, and ASVD across all four model sizes. Moreover, both FWSVD and ASVD demand excessive memory resources, causing out of memory (OOM) when compressing LLMs at larger scales even on an A100 GPU due to memory-intensive operations for estimating the importance of weight matrices. In contrast, SVD-LLM does not involve such estimation operations and thus avoids OOM.
@@ 4.2 Compression Speed Evaluation
Besides compression performance, we also evaluate the compression speed of SVD-LLM and the baselines. Specifically, we measured the GPU hours used for SVD-LLM and ASVD when compressing LLaMA-7B under the 20% compression ratio on an A100 GPU. The results are shown in Section 4.2. As shown, ASVD takes about $5.5$ hours whereas SVD-LLM completes the compression process in $15$ minutes, which is $32$ times faster. When breaking down the time, most of the time consumed by ASVD is dedicated to calculating the compression ratio of each weight matrix based on its estimated importance through a search process. In contrast, SVD-LLM maintains a consistent compression ratio among all weight matrices and thus gets rid of the time-consuming search process.
Table 4: Compression time of SVD-LLM and ASVD on LLaMA-7B under 20% compression ratio. The relative speedup is marked in green color inside bracket.
@@
White Update Total Normalize Search Total
@@
Table 5: Performance of LLaMA-7B compressed by SVD-LLM under 20% and 30% compression ratios using calibration data randomly sampled from WikiText-2 (by default in our paper) and C4. The performance on WikiText-2, PTB, and C4 is reported by perplexity ($\downarrow$), while the performance on OpenbookQA and HellaSwag are reported by accuracy ($\uparrow$). The relative performance drop (gain) for data sampled from C4 compared to that sampled from WikiText-2 is marked in red (green) color inside bracket.
Ratio WikiText-2 PTB C4 Openb. HellaS.
Calibration data sampled from WikiText-2 (seed=3)
20% 7.94 16.22 15.84 0.22 0.43
30% 9.56 26.39 25.11 0.20 0.37
Calibration data sampled from C4 (seed=3)
20% 8.62($\uparrow$9%) 14.63($\downarrow$9%) 13.77($\downarrow$13%) 0.23($\uparrow$5%) 0.43
30% 10.67($\uparrow$12%) 25.07($\downarrow$5%) 22.37($\downarrow$10%) 0.22($\uparrow$9%) 0.42($\uparrow$14%)
4.3 Ablation Study
Table 6: Perplexity ($\downarrow$) of compressed LLaMA-7B on WikiText-2. SVD-LLM (W) denotes the version of SVD-LLM with truncation-aware data whitening only; SVD-LLM (U) denotes the version of SVD-LLM with layer-wise closed-form update only; SVD-LLM (W+U) denotes the version of SVD-LLM with both truncation-aware data whitening and layer-wise closed-form update. The relative performance gain compared to ASVD is marked in green color.